OpenStudy (anonymous):

Find the variance and standard deviation. {3, 3, 4, 5, 5}

OpenStudy (anonymous):

mean = 20 / 5 = 4 3 - mean = -1 3-mean = -1 4 - mean = 0 5 - mean = 1 5 - mean = 1 sigma = 2 variance = sigma (n-mean)^2 / 5 = 4/5 = 0.8 std dev = sqrt 0.8 = 0.89

OpenStudy (anonymous):

THANK you SO much!!! :).

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

plz click on the good answer box to close the question - thanx

OpenStudy (anonymous):

I can't find the button :/.

OpenStudy (anonymous):

on the right hand side of the row where my name is - a blue box

OpenStudy (anonymous):

correction sigma squared = 1 + 1+ 1+ 1 = 4 (sum of the squares of each (number-mean) variance = 0.8 std dev = 0.89 sorry!